Solve and check 3x^2+2x=16
Answer by stanbon(57228)

Solve and check 3x^2+2x=16
------
3x^2+2x-16 = 0
3x^2+8x-6x-16 = 0
x(3x+8)-2(3x+8) = 0
(3x+8)(x-2) = 0
x = 2 or x = -8/3
==================
Cheers,
Stan H.